OpenStudy (anonymous):

Anyone Know how to do this kind of work that will help me through these problems would be really helpful!!!!!!!! 7( x - 1 ) ≥ -14

OpenStudy (anonymous):

divide by 7 get \[x-1\geq -2\] then add 1 to both sides to get \[x\geq -1\]

OpenStudy (saifoo.khan):

\[x-1 \ge - \frac{14}{7}\] \[x-1 \ge - 2\] Add 1 on both sides, \[x \ge -2 +1\] \[x \ge Answer\]

OpenStudy (anonymous):

-3( x - 4 ) - 2( x + 4) ≤ 9

OpenStudy (anonymous):

7x-7>=-14; 7x>=-14+7; x>=-7/7; x>=-1

OpenStudy (anonymous):

-3( x - 4 ) - 2( x + 4) ≤ 9

OpenStudy (anonymous):

\[x\geq -1\] as before

OpenStudy (anonymous):

5( x + 1) - 3( x - 1) > -9

OpenStudy (anonymous):

\[x\geq -\frac{17}{2}\]

OpenStudy (anonymous):

-3x+12-2x-8=<-9; -5x+4=<-9; -5x=<-4-9; -5x=<-13; (x-1) x>=13/5

OpenStudy (anonymous):

4(x + 3) > 6(x - 5)

OpenStudy (anonymous):

4x+12>6x-30; 4x-6x>-30-12; -2x>-42 (x-1) x<42/2; x<21

OpenStudy (anonymous):

x -1 >= -2 ->x >= -1

OpenStudy (anonymous):

3x + 5 < 5

OpenStudy (anonymous):

In general you want to approach inequalities the same way that you would a simple equality however when you multiply or divide by a negative then you must change the direction of the inequality. i.e.) -x > 5 becomes x < 5.

OpenStudy (anonymous):

ok for 3x+ 5< 5 I got x ≥ -3 is that right ?

OpenStudy (anonymous):

sorry -5(x + 3) ≥ 0 I got x ≥ -3